A farmer orders 3 3/4 bales of hay. He spreads 2/5 of the delivery around the stables and stores the rest. How many baes of hay are stored?
Answer by rfer(16322)

3 3/4=15/4
15/4*3/5=45/20=2 5/20=2 1/4 bales stored